OpenStudy (anonymous):

tan(pi/8) Evaluate using half angle identities? Show work please?

OpenStudy (anonymous):

\[ \tan \frac{x}{2} = \frac{\sin x}{1+\cos x}\] Let \[x = \frac{\pi}{4} \implies \frac{x}{2} = \frac{\pi}{8} \] and note: \[\sin \frac{\pi}{4} = \cos \frac{\pi}{4} = \frac{1}{\sqrt{2}} \]